# Discovering Mixtures of Structural Causal Models from Time Series Data

### v2

@@ 6 Conclusion and Discussion
In this work, we examined the problem of discovering mixtures of structural causal models from time series data. This problem has far-reaching applications in climate, finance, and healthcare, among other fields, since multimodal and heterogeneous data is ubiquitous in practice. We proposed an end-to-end variational inference method to learn both the underlying SCMs and the mixture component membership of each sample. We discussed the structural identifiability of mixtures of causal models and demonstrated the empirical efficacy of our method on both synthetic and real-world heterogeneous datasets. Future work could tackle data with latent confounders and non-stationarity in time.
7 Acknowledgement
This work was supported in part by the U.S. Army Research Office under Army-ECASE award W911NF-07-R-0003-03, the U.S. Department Of Energy, Office of Science, IARPA HAYSTAC Program, NSF Grants #2205093, #2146343, and #2134274.
@@ A.1 ELBO derivation
Figure 7: The assumed data generation model. First, the mixture index $Z^{(n)}$ is drawn from a $K$-way categorical distribution ($Z^{(n)}\sim\text{Cat}(K),Z^{(n)}\in\left\{1,\ldots,K\right\}$), and a causal model is drawn from the corresponding mixture component distribution $\mathcal{M}\sim p\left(\mathcal{M}_{Z^{(n)}}\right)$. A sample $X^{(n)}$ is then drawn in accordance with the chosen causal model $\mathcal{M}$.
@@ A.2 Theoretical assumptions
In this section, we list out the theoretical assumptions used in Rhino [Gong et al., 2022]. Our model also operates under similar assumptions, since we implement the component SCMs as Rhino models.
Assumption 1 (Causal Stationarity). [Runge, 2018] The time series $X$ with a graph $G$ is called causally stationary over a time index set $\mathcal{T}$ if and only if for all links $X^{i}_{t-\tau}\rightarrow X_{t}^{j}$ in the graph
@@
Assumption 2 (Causal Markov Property). [Peters et al., 2017] Given a DAG $G$ and a probability distribution $p$, $p$ is said to satisfy the causal Markov property, if it factorizes according to $G$, i.e. $p(x)=\prod_{i=1}^{D}p\left(x_{i}\mid\text{Pa}^{i}_{G}(x_{i})\right)$. In other words, each variable is independent of its non-descendent given its parents.
@@ A.3 Identifiability of the mixture of causal models
Here, we quote a result from Yakowitz and Spragins [1968] that established a necessary and sufficient condition for the identifiability of finite mixtures of multivariate distributions.
Theorem A (Identifiability of finite mixtures of distributions [Yakowitz and Spragins, 1968]).
@@
Proposition A (Identifiability of mixture of multivariate Gaussian distributions [Yakowitz and Spragins, 1968]).
@@
A linear SCM with equal variance Gaussian additive noise is known to be identifiable [Peters and Bühlmann, 2014]. Thus, from Proposition A, we have that the finite mixture is identifiable, as long as the parameters of the resultant Gaussian distributions are distinct, as required by the condition in equation 8.
Conversely, if condition equation 8 does not hold, then $\exists k\neq j$ such that
@@
=$\mathcal{F}=\left\{\mathcal{L}_{\mathcal{M}^{(k)}}:\mathcal{M}^{(k)}\text{ is specified by the equations }\mathbf{X}_{t}=\mathbf{W}^{(k)}\mathbf{X}_{t}+\sum_{\tau=1}^{L}\mathbf{A}^{(k)}_{\tau}\mathbf{X}_{t-\tau}+\varepsilon^{(k)},\varepsilon^{(k)}\sim\mathcal{N}\left(0,\sigma^{2}\mathbf{I}\right),1\leq k\leq K\right\}$
@@
[…]
@@
[…]
@@
Note that $a_{k}$ refers to any point in the support of the mixture distribution such that the condition equation *3 is satisfied. It does not constitute a ’sample’ from the $k^{\text{th}}$ SCM in the conventional sense of being randomly drawn from the SCM. Instead, it can be intentionally chosen to meet the specified condition.
@@ B.2 S&P100
Setup. We provide more details on the setup for the experiment with the S&P 100 dataset. We used grid search to iterate over multiple values for the sparsity term $\lambda$ in equation equation 18, number of graphs $K$, and 5 random seeds. We picked the setting that yielded the lowest validation loss and reported results using the inferred causal graphs. We aggregate the temporal adjacency matrix following the procedure described in Appendix D.5.
@@ B.3 Netsim
Results. In this setup, we observe that MCD-Nonlinear and MCD-Linear are outperformed by the baselines ${}^{+}$ and Rhino, even though they only predict one graph for the entire dataset. This is attributed to the similarity among the various underlying graphs in the Netsim dataset and the strong imbalance in the data. Our model faces sample complexity issues because it learns multiple causal graphs, whereas other methods perform reasonably well by predicting only one. This highlights the idea that learning a mixture model is only beneficial when the underlying SCMs differ from one another significantly. In such a scenario, the benefits of learning multiple graphs outweigh the drawbacks of limited samples per model. This explanation is also supported by the observation that ${}^{+}$ (grouped) achieves lower performance than its single graph counterpart. Further, MCD-Nonlinear and MCD-Linear achieve relatively low clustering accuracy of $35.2\pm 6.6\%$ and $35.4\pm 5.2\%$, due to the inherent similarities in the underlying SCMs.
@@ B.4 Ablation studies
Effect of number of samples per component. We investigate the effect of a decreasing number of samples per mixture component on the performance of MCD, as the number of ground truth SCMs $K^{\ast}$ increases. We consider synthetic nonlinear data of dimension $D=10$ with $K^{\ast}=40,60,80,100$ ground truth graphs in addition to the settings discussed in the main paper. We run MCD-Nonlinear, Rhino and ${}^{+}$ on $N=1000$ samples generated from $K^{\ast}$ SCMs for increasing values of $K^{\ast}$, with $K=2K^{\ast}$. The results are presented in Figure 13. MCD-Nonlinear suffers a gradual decrease in model performance, with roughly a $40\%$ decrease in F1 and $22\%$ decrease in AUROC from $K^{\ast}=1$ to $K^{\ast}=100$. Meanwhile, the performance of Rhino falls off more drastically and becomes equivalent to random guessing for large $K^{\ast}$. The performance of ${}^{+}$ (grouped) also decreases quite rapidly with the increase in $K^{\ast}$.
@@
Robustness of MCD to the misspecification of number of models. We examine the performance of MCD when the number of mixture components $K$ is misspecified, and does not equal the true number of underlying components $K^{\ast}$. Figure 15 shows the performance of MCD-Nonlinear as a function of $K$ on the nonlinear synthetic dataset with dimensionality $D=10$ and ground truth number of graphs $K^{\ast}=10$. We note that when the number of models is underspecified, our model performs poorly as expected since it cannot fully explain all the modes in the data. Surprisingly, the performance increases with increasing $K$. The clustering accuracy and performance metrics show high standard deviation when $K$ is set to the true number of mixture components $K^{\ast}=10$. While some runs achieve high clustering accuracy, others tend to saturate at a suboptimal grouping when $K=K^{\ast}$. On the other hand, when $K>K^{\ast}$, the additional SCMs are used as ‘buffers’, and the correct grouping is learned during the later epochs as the SCMs are inferred more accurately. This phenomenon is further explored in Appendix B.6
@@ B.7 Netsim visualization
Figure 18 shows a visualization of a heatmap of the predictions for the Netsim-mixture dataset. The 3 ground truth adjacency matrices and the top-3 discovered adjacency matrices, ranked by the frequency of prediction, are shown. All 3 matrices achieve a high AUROC score, even though the poor calibration of scores results in the prediction of many spurious edges.
@@ Appendix C Implementation details
In this section, we elaborate more on how we model the terms in equation equation 1 for MCD-Linear and MCD-Nonlinear.
@@
MCD-Nonlinear. We use the Rhino model [Gong et al., 2022] to model each of the $K$ SCMs. The functions $f_{k}$ are modeled as follows:
@@
The first term is a sparsity prior and $h\left(\left(\mathcal{G}_{k}\right)_{0}\right)$ is the acyclicity constraint from [Zheng et al., 2018].
@@ D.1 Synthetic datasets setup
This section provides more details about how we set up and run experiments using MCD on synthetic datasets. We set the number of mixture components $K$ to twice that of true graphs (i.e., $K=2K^{\ast}$) to showcase its robustness against over-specification of the number of components. We set a uniform prior for the mixing probabilities $p\left(Z^{(n)}\right)$, i.e. $p(Z^{(n)}=k)=\frac{1}{K}\,\forall k\in\left\{1,\ldots,K\right\}$. Our implementation of the likelihood function for Rhino-g on the synthetic datasets matches the type of causal relationships modeled, i.e., we use the linear model described in Equation equation 5 on the linear dataset and the nonlinear variant described in Equation equation 6 for the nonlinear datasets.
@@
We generate two separate sets of synthetic datasets: a linear dataset with independent Gaussian noise and a nonlinear dataset with history-dependent noise modeled using conditional splines [Durkan et al., 2019]. We generate a pool of $K^{\ast}$ random graphs (specifically, Erdős-Rényi graphs) and treat them as ground-truth causal graphs. To generate a sample $X^{(n)}$, we assign it to a graph by drawing $Z^{(n)}\sim\text{Categorical}(K^{\ast})$, and use the corresponding graph $\mathcal{G}_{Z^{(n)}}$ from this pool to model relationships between the variables.
@@
where $f^{d}_{k}$ are randomly initialized multi-layer perceptrons (MLPs), and the random noise $\epsilon_{t}^{d}$ is generated using history-conditioned quadratic spline flow functions [Durkan et al., 2019].
@@ D.3 Hyperparameter details
Since MCD uses the acyclicity constraint from Zheng et al. [2018], we use an augmented Lagrangian training procedure to ensure that our model produces DAGs. We closely follow the implementation of the procedure from Geffner et al. [2022], Gong et al. [2022] with one exception: we modify the convergence criteria by increasing the number of outer steps for which the DAG penalty needs to be lower than a threshold (set to $10^{-8}$). This modification enables the model to train for longer and prevents premature stopping. In the interest of fairness, we make this change to both MCD and Rhino.
We used the rational spline flow model described in [Durkan et al., 2019]. We use the quadratic or linear rational spline flow model in all our experiments, both with 8 bins. The MLPs $\ell$ and $\xi$ have 2 hidden layers each, with hidden dimensions set to $\max\left(4D,e,64\right)$ with LeakyReLU activation functions, where $e$ is the embedding dimension. We also use layer normalization and skip connections. The temperature for sampling the adjacency matrix from $q_{\phi}\left(\mathcal{M}_{1:K}\right)$ using the Gumbel Softmax distribution was set to 0.25, and the temperature $\tau_{r}$ for sampling from the mixing rates variational distribution was set to 1. We used the same hyperparameters for both MCD-Nonlinear and MCD-Linear. Table 3 summarizes the hyperparameters used for training.
@@
Sparsity factor $\lambda$ 5 25 10 10 25
@@
D.4 Post-processing the output of ${}^{+}$
${}^{+}$ produces Markov equivalence classes rather than fully oriented causal graphs for the instantaneous adjacency matrix. To make its outputs comparable, we post-process the resultant edges. We follow the setup in Gong et al. [2022] and enumerate up to 3000 DAGs for the instantaneous matrix. We ignore the edges (i.e., set the corresponding entries in the adjacency matrix to 0) whose orientations are undecided. We compare all outputs against the ground truth during the evaluation and return the average metric across all enumerations.

### v3

@@ 6 Conclusion and Discussion
In this work, we examine the problem of discovering mixtures of structural causal models from time series data. This problem has far-reaching applications in climate, finance, and healthcare, among other fields, since multimodal and heterogeneous data is ubiquitous in practice. We propose MCD, an end-to-end variational inference method, to learn both the underlying SCMs and the mixture component membership of each sample. We demonstrate the empirical efficacy of our method on both synthetic and real-world heterogeneous datasets. We conduct ablation studies on synthetic datasets to investigate MCD’s behavior with varying numbers of causal graphs, its robustness to misspecification of the number of graphs, and the impact of similarity among the causal graphs associated with each mixture component. In addition, we discuss the structural identifiability of mixtures of causal models. Future work could tackle data with latent confounders and non-stationarity in time.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Acknowledgement
This work was supported in part by the U.S. Army Research Office under Army-ECASE award W911NF-07-R-0003-03, the U.S. Department Of Energy, Office of Science, IARPA HAYSTAC Program, NSF Grants SCALE MoDL-2134209, CCF-2112665 (TILOS), #2205093, #2146343, #2134274, CDC-RFA-FT-23-0069 and DARPA AIE FoundSci.
@@ A.1 ELBO derivation
Figure 8: The assumed data generation model. First, the mixture index $Z^{(n)}$ is drawn from a $K$-way categorical distribution ($Z^{(n)}\sim\text{Cat}(K),Z^{(n)}\in\left\{1,\ldots,K\right\}$), and a causal model is drawn from the corresponding mixture component distribution $\mathcal{M}\sim p\left(\mathcal{M}_{Z^{(n)}}\right)$. A sample $X^{(n)}$ is then drawn in according to the chosen causal model $\mathcal{M}$.
@@ A.2 Theoretical assumptions
In this section, we list out the theoretical assumptions used in Rhino (Gong et al., 2022). Our model also operates under similar assumptions, since we implement the component SCMs as Rhino models.
Assumption 1 (Causal Stationarity). (Runge, 2018) The time series $X$ with a graph $G$ is called causally stationary over a time index set $\mathcal{T}$ if and only if for all links $X^{i}_{t-\tau}\rightarrow X_{t}^{j}$ in the graph
@@
Assumption 2 (Causal Markov Property). (Peters et al., 2017) Given a DAG $G$ and a probability distribution $p$, $p$ is said to satisfy the causal Markov property, if it factorizes according to $G$, i.e. $p(x)=\prod_{i=1}^{D}p\left(x_{i}\mid\text{Pa}^{i}_{G}(x_{i})\right)$. In other words, each variable is independent of its non-descendent given its parents.
@@ A.3 Identifiability of the mixture of causal models
Here, we quote a result from Yakowitz & Spragins (1968) that established a necessary and sufficient condition for the identifiability of finite mixtures of multivariate distributions.
Theorem A (Identifiability of finite mixtures of distributions (Yakowitz & Spragins, 1968)).
@@
Proposition A (Identifiability of mixture of multivariate Gaussian distributions (Yakowitz & Spragins, 1968)).
@@
A linear SCM with equal variance Gaussian additive noise is known to be identifiable (Peters & Bühlmann, 2014). Thus, from Proposition A, we have that the finite mixture is identifiable, as long as the parameters of the resultant Gaussian distributions are distinct, as required by the condition in (8).
Conversely, if condition (8) does not hold, then $\exists k\neq j$ such that
@@
=$\mathcal{F}=\Biggl{\{}\mathcal{L}_{\mathcal{M}^{(k)}}:\mathcal{M}^{(k)}\text{ is specified by the equations }\mathbf{X}_{t}=\mathbf{W}^{(k)}\mathbf{X}_{t}$ $+\sum_{\tau=1}^{L}\mathbf{A}^{(k)}_{\tau}\mathbf{X}_{t-\tau}+\varepsilon^{(k)},$
$\varepsilon^{(k)}\sim\mathcal{N}\left(0,\sigma^{2}\mathbf{I}\right),1\leq k\leq K\Biggr{\}}$
@@
where $a_{k}$ is as defined in Condition (($\right)$ ). Denote $\mathcal{L}_{\mathcal{M}_{j}}(a_{k})=\beta_{kj}$. Then Equation (10) can be written as:
@@
Note that $\bm{\alpha}=0$ if and only if $\bm{\beta}$ is full rank. We now show that Condition (($\right)$ ) implies that $\bm{\beta}$ is strictly diagonally dominant and hence full rank. Note that Condition (($\right)$ ) can be equivalently written as:
@@
Note that $a_{k}$ refers to any point in the support of the mixture distribution such that the condition (*3) is satisfied. It does not constitute a ’sample’ from the $k^{\text{th}}$ SCM in the conventional sense of being randomly drawn from the SCM. Instead, it can be intentionally chosen to meet the specified condition.
@@ B.2 S&P100
Setup. We provide more details on the setup for the experiment with the S&P 100 dataset. We used grid search to iterate over multiple values for the sparsity term $\lambda$ in equation (18), number of graphs $K$, and 5 random seeds. We picked the setting that yielded the lowest validation loss and reported results using the inferred causal graphs. We aggregate the temporal adjacency matrix following the procedure described in Appendix D.5.
@@ B.3 Netsim
Results. In this setup, we observe that MCD-Nonlinear and MCD-Linear are outperformed by the baselines PCMCI+ and Rhino, even though they only predict one graph for the entire dataset. This is attributed to the similarity among the various underlying graphs in the Netsim dataset and the strong imbalance in the data. Our model faces sample complexity issues because it learns multiple causal graphs, whereas other methods perform reasonably well by predicting only one. This highlights the idea that learning a mixture model is only beneficial when the underlying SCMs differ from one another significantly. In such a scenario, the benefits of learning multiple graphs outweigh the drawbacks of limited samples per model. This explanation is also supported by the observation that PCMCI+ (grouped) achieves lower performance than its single graph counterpart. Further, MCD-Nonlinear and MCD-Linear achieve relatively low clustering accuracy of $35.2\pm 6.6\%$ and $35.4\pm 5.2\%$, due to the inherent similarities in the underlying SCMs.
@@ B.4 Ablation studies
Effect of number of samples per component. We investigate the effect of a decreasing number of samples per mixture component on the performance of MCD, as the number of ground truth SCMs $K^{\ast}$ increases. We consider synthetic nonlinear data of dimension $D=10$ with $K^{\ast}=40,60,80,100$ ground truth graphs in addition to the settings discussed in the main paper. We run MCD-Nonlinear, Rhino and PCMCI+ on $N=1000$ samples generated from $K^{\ast}$ SCMs for increasing values of $K^{\ast}$, with $K=2K^{\ast}$. The results are presented in Figure 14. MCD-Nonlinear suffers a gradual decrease in model performance, with roughly a $40\%$ decrease in F1 and $22\%$ decrease in AUROC from $K^{\ast}=1$ to $K^{\ast}=100$. Meanwhile, the performance of Rhino falls off more drastically and becomes equivalent to random guessing for large $K^{\ast}$. The performance of PCMCI+ (grouped) also decreases quite rapidly with the increase in $K^{\ast}$.
@@
Effect of similarity of the graphs on performance
Figure 16: Causal discovery performance and clustering accuracy of MCD on synthetic datasets with similar graphs. MCD achieves good causal discovery performance and clustering accuracy even when $-\log p$ is high, i.e., the graphs are similar.
We examine the performance of MCD when the causal graphs are similar to each other. The dataset setup is as follows: we first generate a random ER graph. We then perturb each edge with a probability $p$, i.e. flip the entry in the adjacency matrix from 0 to 1, and vice versa with probability $p$. We check if the resulting graph is a DAG. If yes, we add it to the pool of generating graphs. We repeat this procedure until we obtain $K^{*}$ DAGs. We then generate the synthetic dataset ($N=1000$ samples) with these $K^{*}$ DAGs using randomly generated MLPs and spline functions. We run MCD-Nonlinear with $K=2K^{*}$ on the resulting datasets with $D=10$ and $K=5,10$. We set $p=0.005,0.008,0.01,0.05,0.1$, corresponding to varying levels of similarity of the underlying graphs. We also attach the pair-wise statistics for the resulting graphs in Table 4. The results are reported in Figure 16
The results indicate that MCD can achieve good clustering accuracy and causal discovery performance even when the constituent causal graphs are similar. In the $K^{*}=5$ case, we notice that the clustering accuracy remains high for all considered settings, and the F1 score decreases slightly for the higher values of $-\log p$ when the causal graphs are similar. For $K^{*}=10$, the clustering accuracy is considerably lower for $p=0.005$, but nevertheless the F1 score remains high. On the other hand, PCMCI+-g suffers a gradual performance drop as the graphs become similar, while the single graph baselines Rhino and PCMCI+-s have slightly better performance when the constituent causal graphs are more similar.
@@ B.7 Netsim visualization
Figure 19 shows a visualization of a heatmap of the predictions for the Netsim-mixture dataset. The 3 ground truth adjacency matrices and the top-3 discovered adjacency matrices, ranked by the prediction frequency, are shown. All 3 matrices achieve a high AUROC score, even though the poor calibration of scores results in the prediction of many spurious edges.
@@ Appendix C Implementation details
In this section, we elaborate more on how we model the terms in equation (1) for MCD-Linear and MCD-Nonlinear.
@@
MCD-Nonlinear. We use the Rhino model (Gong et al., 2022) to model each of the $K$ SCMs. The functions $f_{k}$ are modeled as follows:
@@
The first term is a sparsity prior and $h\left(\left(\mathcal{G}_{k}\right)_{0}\right)$ is the acyclicity constraint from (Zheng et al., 2018).
@@ D.1 Synthetic datasets setup
This section provides more details about how we set up and run experiments using MCD on synthetic datasets. We set the number of mixture components $K$ to twice that of true graphs (i.e., $K=2K^{\ast}$) to showcase its robustness against over-specification of the number of components. We set a uniform prior for the mixing probabilities $p\left(Z^{(n)}\right)$, i.e. $p(Z^{(n)}=k)=\frac{1}{K}\,\forall k\in\left\{1,\ldots,K\right\}$. Our implementation of the likelihood function for Rhino-g on the synthetic datasets matches the type of causal relationships modeled, i.e., we use the linear model described in Equation (5) on the linear dataset and the nonlinear variant described in Equation (6) for the nonlinear datasets.
@@
We generate two separate sets of synthetic datasets: a linear dataset with independent Gaussian noise and a nonlinear dataset with history-dependent noise modeled using conditional splines (Durkan et al., 2019). We generate a pool of $K^{\ast}$ random graphs (specifically, Erdős-Rényi graphs) and treat them as ground-truth causal graphs. To generate a sample $X^{(n)}$, we assign it to a graph by drawing $Z^{(n)}\sim\text{Categorical}(K^{\ast})$, and use the corresponding graph $\mathcal{G}_{Z^{(n)}}$ from this pool to model relationships between the variables.
@@
where $f^{d}_{k}$ are randomly initialized multi-layer perceptrons (MLPs), and the random noise $\epsilon_{t}^{d}$ is generated using history-conditioned quadratic spline flow functions (Durkan et al., 2019).
@@ D.3 Hyperparameter details
Since MCD uses the acyclicity constraint from Zheng et al. (2018), we use an augmented Lagrangian training procedure to ensure that our model produces DAGs. We closely follow the implementation of the procedure from Geffner et al. (2022); Gong et al. (2022) with one exception: we modify the convergence criteria by increasing the number of outer steps for which the DAG penalty needs to be lower than a threshold (set to $10^{-8}$). This modification enables the model to train for longer and prevents premature stopping. In the interest of fairness, we make this change to both MCD and Rhino.
We used the rational spline flow model described in (Durkan et al., 2019). We use the quadratic or linear rational spline flow model in all our experiments, both with 8 bins. The MLPs $\ell$ and $\Xi$ have 2 hidden layers each, with hidden dimensions set to $\max\left(4D,e,64\right)$ with LeakyReLU activation functions, where $e$ is the embedding dimension. We also use layer normalization and skip connections. The temperature for sampling the adjacency matrix from $q_{\phi}\left(\mathcal{M}_{1:K}\right)$ using the Gumbel Softmax distribution was set to 0.25, and the temperature $\tau_{r}$ for sampling from the mixing rates variational distribution was set to 1. We used the same hyperparameters for both MCD-Nonlinear and MCD-Linear. Table 3 summarizes the hyperparameters used for training.
@@
Sparsity factor $\lambda$ 5 25 10 20 25
@@
D.4 Post-processing the output of PCMCI+
PCMCI+ produces Markov equivalence classes rather than fully oriented causal graphs for the instantaneous adjacency matrix. To make its outputs comparable, we post-process the resultant edges. We follow the setup in Gong et al. (2022) and enumerate up to 3000 DAGs for the instantaneous matrix. We ignore the edges (i.e., set the corresponding entries in the adjacency matrix to 0) whose orientations are undecided. We compare all outputs against the ground truth during the evaluation and return the average metric across all enumerations.